### v1

# VerifyThisBench: Generating Code, Specifications, and Proofs All at Once

Abstract
Large language models (LLMs) have demonstrated remarkable progress in code generation, but many existing benchmarks are approaching saturation and offer little guarantee on the trustworthiness of the generated programs, offering limited insight into deeper reasoning capabilities. We introduce VerifyThisBench, a new benchmark designed to evaluate LLMs on end-to-end program verification tasks that require interpreting natural language problem descriptions, formulating formal specifications, generating code, and constructing correctness proofs. Our evaluation reveals that even state-of-the-art (SOTA) models, such as o3-mini, achieve a pass rate of less than 4%, with many outputs failing to compile. To reduce task complexity, we further propose VerifyThisBenchXS, a variant in which partial implementations or proofs are provided. We systematically assess SOTA models on both benchmarks, uncovering key strengths and limitations in their formal reasoning and verification capabilities.
Keywords LLM  $\cdot$ Formal Verification  $\cdot$ Program Synthesis  $\cdot$ Machine Learning for Formal Methods
1 Introduction
Large Language Models (LLMs) have unequivocally revolutionized the landscape of automated code generation. Models like GPT-4o [33], Google Gemini [15], Anthropic Claude [1], and GitHub Copilot [14] excel at generating functional code snippets and translating between languages. These capabilities are now integrated into AI-powered IDEs, such as Cursor AI, to support large-scale software development. This proficiency had led to increasing needs on established benchmarks, as early as HumanEval [6] and MBPP [3], to reflect the capability of each LLM tool. However, this rapid progress raises critical questions about the trustworthiness and reliability of the generated artifacts. Many existing benchmarks, while useful for gauging functional correctness through test suites, are approaching saturation [21, 13, 16, 45] and inherently offer limited guarantees regarding the deeper aspects of program correctness. Test cases, by their nature, can demonstrate the presence of bugs, but cannot prove their absence [11], leaving a significant gap in assessing the formal robustness and true reasoning capabilities of these powerful models.
Reliable software must go beyond passing tests to be trustworthy, precisely follow specifications, and even self-validate. Formal verification offers the most rigorous approach to achieving these guarantees. This paradigm involves providing mathematical proof that a program adheres to a formal specification, thereby guaranteeing critical properties such as functional correctness, liveness (ensuring the program eventually does something good), and safety (ensuring the program never does something bad) [18]. Modern program verification infrastructures, such as Dafny [27], Frama-C [22], Verus [26], Isabelle/HOL [32], and Lean [8], coupled with powerful automated theorem provers and SMT solvers like Z3 [7] and CVC5 [4], have significantly streamlined the process of writing and checking such verified software. These tools allow developers to express complex specifications and then automatically or semi-automatically verify that the implementation meets these specifications.
Although researchers have developed multiple benchmarks to assess LLMs on formal verification subtasks [20, 5, 35, 12], none evaluate end-to-end program verification starting from a natural-language description. Instead, existing suites either require verifying or synthesizing small programs against a formal specification or focus on aiding proof completion by suggesting individual verification steps. Consequently, even though state-of-the-art LLMs have been reported to solve up to 97.8% of these benchmark tasks [44], those numbers do not reflect their true capability for end-to-end program verification.
To bridge this gap and rigorously evaluate the capabilities of LLMs in this demanding domain, we introduce VerifyThisBench, a novel benchmark designed to assess end-to-end program verification, as shown in Figure 1. Inspired by the annual VerifyThis Challenge [43], where human contestants devise implementations and accompanying formal proofs in verification-aware languages, VerifyThisBench tasks LLMs with interpreting natural language problem descriptions, formulating formal specifications, generating the corresponding code, and constructing machine-checkable correctness proofs—all at once.
Our evaluation using VerifyThisBench reveals that even state-of-the-art (SOTA) models, such as o3-mini, achieve a zero-shot pass rate of 3.62% on this end-to-end task, with a significant number of outputs failing even to compile, and only reach a pass rate of 9.37% after five rounds of feedback. These results underscore the profound challenge this domain presents. To dissect these challenges further and explore capabilities in a more guided setting, we also propose VerifyThisBenchXS, a variant where partial implementations or proofs are provided, and the LLM’s task is to complete the missing components.
This paper makes the following key contributions:
  • VerifyThisBench: We present VerifyThisBench, a new benchmark suite for evaluating the ability of LLMs to generate fully verified programs (code, specifications, and proofs) from natural language descriptions.
  • Relaxed VerifyThisBench: We introduce VerifyThisBenchXS, a relaxed version of the VerifyThisBench, to assess LLM performance when provided with partial artifacts and tasked with completing them.
  • Unified Environment: We provide a unified evaluation environment that integrates seven verification tools and an automated pipeline, enabling consistent and scalable benchmarking across diverse formal verification tasks.
  • SOTA LLM Evaluation: We conduct a systematic evaluation of nine SOTA LLMs on both benchmarks, revealing current capabilities and significant limitations. Our analysis includes performance breakdowns across tools, attempt-based comparisons, model-specific strengths, self-assessed coherence, and the impact of partial guidance, providing a comprehensive understanding of model behavior in formal verification tasks.
    Figure 1: Evaluation workflow of VerifyThisBench and its relaxed settings.
2 Background & Related Work
2.1 Unverified Code Synthesis Benchmarks
Recent benchmarks for code generation include APPS [17], HumanEval [6], MBPP [3], CodeContests [28], DS-1000[25], SWEBench [19], and EvalPlus [29], among others. These benchmarks present programming tasks, often sourced from online competitions or community platforms, and evaluate models based on whether generated solutions pass a set of input-output test cases. While effective for measuring functional correctness, they do not involve formal specifications, proofs, or specification synthesis.
In contrast, VerifyThisBench requires models to go beyond functional testing: they must extract a formal specification from a natural-language description, generate code in a verification-aware language, and produce a proof that passes formal verification. This makes VerifyThisBench a substantially more rigorous and comprehensive benchmark than traditional synthesis tasks.
2.2 Program Verification Benchmarks
Benchmarks in formal verification include SV-COMP [37], SyGuS [38], and Code2Inv [36]. SV-COMP and Code2Inv focus solely on verification task with no code generation involved. Specifically, the former contains large-scale C benchmarks with fixed safety properties and the latter targets invariant generation over small C-style programs. SyGuS focuses on constraint-based synthesis.
More recent efforts like DafnyBench  [30] and VerusBench  [46] collect verified programs in Dafny and Verus respectively, primarily to train and evaluate ML-based tools for aiding in proof completion and suggesting verification steps, rather than end-to-end program generation from natural language.
These benchmarks evaluate components of the verification pipeline but typically assume a preset formal specification or verification goal. In contrast, VerifyThisBench uses the end-to-end setup to explicitly evaluate the model’s ability in interpreting and encoding natural-language descriptions into provably correct formal programs, a capability not tested in existing benchmarks.
2.3 Formal Methods in Software Verification: A Primer
Formal methods in software verification aim to mathematically prove program correctness against a formal specification—a precise, unambiguous description of what a program should do, often expressed in a logical language. This contrasts with testing, which can only show the presence of bugs for specific inputs. The verification process typically relies on several key components embedded within or alongside the executable program code:
  • Contracts: These formalize the obligations and guarantees of a code segment.
    • Pre-conditions (requires clauses): Properties that must hold true before a function or code block executes for it to behave correctly.
    • Post-conditions (ensures clauses): Properties guaranteed to be true after a function or code block finishes, provided its pre-conditions were met.
  • Intermediate Assertions: Assistive hints are often needed to bridge any reasoning gaps between the pre&post-conditions where the underlying solver cannot automatically address.
  • Loop Invariants: For iterative constructs, loop invariants are crucial properties that hold at the start of a loop, are preserved by each iteration, and, in conjunction with the loop’s termination, help prove the loop’s correctness.
The typical verification flow in systems utilizing these concepts is as follows:
  1. 1.
    Annotation: Developers write code in a verification-aware language (e.g., Dafny [27], Frama-C [22], Verus [26] and annotate it with formal specifications and proof hints, including pre-conditions, post-conditions, assertions, and loop invariants.
  2. 2.
    Generation of Proof Obligations: A tool, often a Verification Condition Generator (VCG), processes the annotated code and its specifications. It translates them into a series of mathematical proof obligations (verification conditions) that, if all true, logically imply the program’s correctness with respect to its specification.
  3. 3.
    Automated Proving: These verification conditions are then fed to backend automated theorem provers, typically Satisfiability Modulo Theories (SMT) solvers like Z3 [7] or CVC5 [4]. These solvers attempt to mathematically prove each obligation.
  4. 4.
    Feedback: The system reports to the developer whether the proofs succeeded or failed. Failures often pinpoint inconsistencies between the code and its specification, or missing/incorrect annotations.
Successfully generating code within this paradigm, as targeted by our VerifyThisBench benchmark, requires an LLM not only to produce the algorithmic implementation but also to understand, formulate, and correctly express these intricate formal specifications and proof structures that enable automated verification.
3 VerifyThisBench Benchmark
VerifyThisBench is inspired by the annual VerifyThis Challenges [43], a competition where participants are tasked with formalizing specifications, implementing solutions, and verifying that the implementations meet the specification. Each challenge is designed to be completed within a 90-minute session and varies in difficulty. Submissions are evaluated based on correctness, completeness, and additional quality criteria such as elegance and the degree of automation. Similarly, in VerifyThisBench, the task is to interpret natural language problem descriptions and implement code and write proofs.
3.1 Benchmark Construction
We collected challenges from each annual competition between 2011 and 2024, documenting their descriptions, pseudo-code, and associated tasks. Tasks are categorized as either implementation (completing an algorithm from pseudo-code) or verification (proving a model or implementation correct against a specification). All tasks are described in natural language. In total, the dataset includes 41 challenges and 154 tasks. The dataset is available at [10].
3.2 Environment
To facilitate evaluation, we provide a unified environment supporting seven verification tools. Five of them, Dafny [27], Why3 [42], VeriFast [41], VerCors [40], and Frama-C [22], are widely used in past VerifyThis competitions. To broaden tool diversity, we additionally include Verus [26] and CBMC [23]. Tool versions and brief descriptions can be found in Appendix D.
3.3 Features of VerifyThisBench
End-to-end verification tasks with natural language problem descriptions: All tasks start with informal, natural language prompts (often with pseudo-code). Models must interpret the intent and formalize it into precise logical specifications. They are required to generate specifications, implementations, and formal proofs in a verification-aware language, ensuring the code passes machine-checkable verification. Example challenge and solution can be found in Appendix E.
Graded difficulty and multi-step challenges: Challenges are drawn from the VerifyThis competition and span a range of difficulties. Many include sequential subtasks, allowing fine-grained assessment of model capability.
Tool diversity: Multiple tools are provided. Models must conform to the syntax and semantics of real-world verification frameworks.
3.4 Relaxation
We observe that most language models fail to generate compilable code when targeting specific formal verification tools. This is often due to the syntactic complexity and precise annotations required by these tools. To decrease the level of difficulty and better assess LLM capabilities under more supportive conditions, we construct a set of relaxed problems derived from past human-written solutions. Specifically, we define three forms of relaxation:
  • Code relaxation: We provide only the function specifications, omitting both the implementation and the proof annotations.
  • Specification relaxation: We provide the implementation and its proof, but remove the function specifications.
  • Proof relaxation: We provide specifications and implementations, but remove loop invariants and other auxiliary annotations needed for verification.
To further diversify the difficulty spectrum, we vary the extent of relaxation. In some instances, we remove all relevant components (e.g., entire specs or proofs), while in others, we retain partial elements or include complete examples as guidance. This enables a more graded evaluation of LLM performance across varying levels of verification support. In total, we create a set of 481 tasks. Specifically, there are 195 fill-implementation task, 90 fill-proof/invariants tasks, and 196 fill-specification tasks. Table 6 in Appendix A shows the statistics of VerifyThisBenchXS. As there are no prior solutions in CBMC and Verus, no tasks were created and no results reported for these tools in the relaxed setting.
4 Experiment Results
4.1 Evaluation Setup and Metrics
We evaluate a diverse set of state-of-the-art (SOTA) language models, including both proprietary and open-source systems. The models include representatives from the OpenAI family (GPT-4o, GPT-4omini, o3-mini, o4-mini) [34], Anthropic (Claude-3.7-Sonnet) [2], Google (Gemini-2.5-Flash), DeepSeek (Deepseek-chat-v3) [9], Meta (Llama3.3-70B-Instruct) [31] and Alibaba (Qwen-2.5-72B-Instruct) [39]. This selection enables a comprehensive comparison across different model architectures and training paradigms. Model versions are provided in appendix C.
4.2 Experiment Design
For both VerifyThisBench and VerifyThisBenchXS, we conduct experiments with iterative refinement based on tool-generated error messages. To evaluate correctness, we pass the generated code to the target verification tool and check whether it compiles and verifies successfully. A task is marked as pass if no errors are returned.
In addition to correctness checking, we introduce a coherence check as a relaxed evaluation metric. Here, the model self-assesses whether its generated code semantically aligns with the original problem intent—an aspect difficult to verify automatically. This helps evaluate how well the specification matches the task description and provides insight into the model’s confidence in its output.
Each task is attempted five times per model. The first attempt uses only the task prompt; the next four incorporate feedback from previous errors. During refinement, the model has access to the full history of its prior attempts and corresponding feedback for the current task, enabling iterative improvement.
In VerifyThisBench, tasks of a challenge are completed sequentially. Only the final attempt from the previous task is carried over to the next, preserving essential context while keeping prompts concise. In contrast, VerifyThisBenchXS tasks have isolated contexts and are completed independently, with no progress carried over between tasks.
To ensure fairness, we use the same prompt across all models and set the temperature to 0.7 when applicable. A timeout of 1 minute is enforced for all experiments on the verifier. The experiments were conducted on a machine with an Intel i7-1360P CPU and 16GB of RAM.
4.3 Overall Pass Rate
Table 1: Overall Pass Rate On VerifyThisBench
Attempt GPT4o GPT4o-mini o3-mini o4-mini Claude Gemini Llama Deepseek Qwen
CBMC zero-shot 8.44% 7.14% 8.44% 1.30% 6.49% 1.95% 7.14% 0.65% 1.30%
refinement 20.13% 19.48% 25.32% 15.58% 22.08% 14.94% 20.13% 22.08% 3.25%
Dafny zero-shot 0 0 4.55% 1.95% 3.90% 0 0 1.30% 0
refinement 1.30% 0.65% 9.74% 10.39% 11.04% 1.30% 2.60% 2.60% 0.65%
Frama-C zero-shot 0 0.65% 0 0 3.90% 0.65% 0 1.95% 0
refinement 7.14% 1.95% 2.60% 3.25% 11.04% 8.44% 0.65% 3.25% 0
VerCors zero-shot 0 1.30% 1.30% 1.95% 0 5.84% 8.44% 1.95% 0
refinement 1.95% 1.30% 1.95% 5.19% 1.30% 16.88% 11.69% 4.55% 3.25%
VeriFast zero-shot 0 0 0 0 0 0 0 0 0
refinement 0 0 0 2.60% 0 0 0.65% 0.65% 0
Verus zero-shot 1.95% 6.49% 10.39% 0.65% 0.65% 0.65% 7.79% 0.65% 0.65%
refinement 12.99% 9.09% 21.43% 8.44% 1.30% 0.65% 17.53% 1.30% 0.65%
Why3 zero-shot 0 0 0.65% 0.65% 1.30% 1.30% 0 0.65% 0
refinement 0 0 4.55% 10.39% 9.09% 5.84% 1.95% 1.95% 0
Overall zero-shot 1.48% 2.23% 3.62% 0.93% 2.32% 1.48% 3.34% 1.02% 0.28%
refinement 6.22% 4.64% 9.37% 7.98% 7.98% 6.86% 7.88% 5.19% 1.11%
Improvement 4.73% 2.41% 5.75% 7.05% 5.66% 5.38% 4.55% 4.17% 0.83%
Table 2: Overall Pass Rate On VerifyThisBenchXS
Attempt GPT4o GPT4o-mini o3-mini o4-mini Claude Gemini Llama Deepseek Qwen
Dafny zero-shot 0 1.35% 2.70% 1.35% 1.35% 1.35% 0 2.70% 1.35%
refinement 17.57% 9.46% 31.08% 37.84% 41.89% 17.57% 8.11% 21.62% 8.11%
Frama-C zero-shot 0 0 1.85% 0 1.85% 1.85% 0 1.85% 0
refinement 0 0 5.56% 18.52% 9.26% 1.85% 0 5.56% 0
VerCors zero-shot 0 0 0 0 0 0 0 0 0
refinement 0 0 0 7.69% 0 0 3.85% 0 0
VeriFast zero-shot 7.58% 4.55% 3.03% 6.06% 6.06% 0 4.55% 12.12% 3.03%
refinement 12.12% 6.06% 4.55% 10.61% 27.27% 0 9.09% 13.64% 6.06%
Why3 zero-shot 0 0 0 0.38% 0.38% 0.77% 0 0.38% 0
refinement 7.66% 2.30% 0.77% 8.81% 3.45% 1.15% 15.71% 18.77% 1.15%
Overall zero-shot 1.04% 0.83% 1.04% 1.25% 1.46% 0.83% 0.62% 2.49% 0.62%
refinement 8.52% 3.53% 6.44% 14.55% 13.10% 3.53% 11.23% 16.01% 2.70%
Improvement 7.48% 2.70% 5.40% 13.30% 11.64% 2.70% 10.61% 13.52% 2.08%
Table 1 presents the performance of the SOTA models on VerifyThisBench. For each verification tool, we report pass rates for the initial zero-shot attempt and after four additional refinement attempts using feedback.
In the first attempt, most models perform poorly, with success rates under 4%. The top performers are o3-mini, Llama, and Claude, indicating that even the strongest models struggle initially. By the fifth attempt, performance improves significantly across all models. o3-mini leads overall, followed by Claude, o4-mini, and Llama. These results highlight the effectiveness of iterative refinement and feedback in enhancing model performance.
Each model exhibits distinct strengths across different verification tools, underscoring that no single model consistently outperforms the rest. For example, o3-mini, the top overall performer, excels especially in CBMC and Verus. On the other hand, Claude shows consistent strength in Dafny and Frama-C. Gemini, while generally average, performs exceptionally well on VerCors. Llama, another open-source model, performs best on Verus. In contrast, Qwen shows consistently low performance across all tools, suggesting limitations in its current proof synthesis capabilities. Further insights into tool-specific performance are discussed in Section 4.6.
Table 2 shows the results on VerifyThisBenchXS. Similarly, at the first attempt, absolute numbers remain low (less than 3%) for all models. At the fifth iteration, Deepseek tops the competition with 16.01%, followed closely by o4-mini (14.55%), Claude (13.10%), and Llama (11.23%). Feedback leads to substantial improvement for most models, achieving relative gains of over 10%.
In conclusion, while few models succeed from scratch, many become competitive when guided by partial context. Open-source models like Deepseek, and Llama outperform many closed-source counterparts, showing strong potential for real-world deployment in assisted formal verification. These results also underscore the importance of combining structural hints, feedback loops, and domain-specific strengths when applying LLMs to formal reasoning tasks.
Key Insights: Average pass rates remain low at 10% on VerifyThisBench and 16% on VerifyThisBenchXS, revealing the challenges formal verification poses even to SOTA LLMs. All models improve with feedback.
4.4 Failure Mode Distribution
Figure 2: zero-shot on VerifyThisBench
Figure 3: refinement on VerifyThisBench
Figure 4: zero-shot on VerifyThisBenchXS
Figure 5: refinement on VerifyThisBenchXS
Figures 5 to 5 show clear improvements in model’s performance when partial solution templates are provided.
Specifically, partial success rates, where the verifier is able to confirm some goals, increase significantly. This suggests that templates or hints help models generate more accurate solutions. Timeout rates, where the program compiles but the verifier fails to complete within the time limit, remain relatively stable. This indicates that models are making meaningful progress toward valid proofs, and the verifier struggles to find counterexamples, implying the generated solutions are closer to correctness.
Compilation errors still dominate but tend to decrease under the relaxed setting for some models, demonstrating that context helps reduce syntax-level mistakes. However, some models like GPT4o-mini and o3-mini exhibit mixed trends, suggesting that while the template helps, the model’s internal understanding and code generation fidelity still vary.
If we relax the metric to consider compilable code rather than fully verified solutions, Claude, GPT-4o, and Deepseek consistently emerge as the top performers across both benchmarks. Notably, Claude generates compilable outputs in nearly 50% of attempts on VerifyThisBenchXS and around 25% on VerifyThisBench in the first attempt alone, highlighting its strong baseline capability even without iterative feedback.
Key Insights: While compilation error dominates in both benchmarks, in the relax setting we observe decreases in such failures and increases of partial correct or compilable solutions, moving model performance closer to usable verification outputs even when full correctness is not achieved.
4.5 Coherence
Table 3: Self-Assessment of Specification Coherence on VerifyThisBench
Attempt GPT4o GPT4o-mini o3-mini o4-mini Claude Gemini Llama Deepseek Qwen
CBMC zero-shot 15.38% 0% 84.62% 0% 80.00% 33.33% 0% 100% 100%
refinement 16.13% 0% 61.54% 12.50% 26.47% 13.64% 3.23% 20.59% 100%
Dafny zero-shot - - 100% 100% 100% - - 100% -
refinement 50.00% 100% 100% 62.50% 76.47% 50.00% 25.00% 100% 0%
Frama-C zero-shot - 100% - - 100% 0% - 100% -
refinement 100% 100% 75.00% 80.00% 70.59% 53.85% 0% 100% -
VerCors zero-shot - 100% 100% 33.33% - 66.67% 69.23% 100% -
refinement 66.67% 100% 100% 62.50% 0% 46.15% 61.11% 85.71% 80.00%
VeriFast zero-shot - - - - - - - - -
refinement - - - 0% - - 0% 100% -
Verus zero-shot 0% 30.00% 100% 0% 0% 0% 8.33% 0% 0%
refinement 0% 28.57% 93.94% 7.69% 0% 0% 3.70% 0% 0%
Why3 zero-shot - - 100% 100% 100% 0% - 100% -
refinement - - 100% 53.33% 35.71% 22.22% 0% 100% -
Average zero-shot 12.50% 25.00% 94.87% 50.00% 88.00% 43.75% 27.78% 90.91% 66.67%
refinement 28.36% 20.00% 82.00% 36.47% 45.35% 34.25% 16.47% 46.43% 75.00%
Table 3 reports each model’s coherence confidence, i.e. whether the model believes its generated specification matches the intended problem requirement. This metric is evaluated across generations passing the verification tool. While passing a formal verifier indicates syntactic and logical correctness, it does not address the alignment problem (i.e., whether the verified implementation perfectly algins with the user-intent expressed in natural language descriptions); hence, coherence offers complementary insight. Notably, except o3-mini and Qwen, models’s confidence is less than 50% on passed solutions.
The results reveal considerable variance across models in their self-assessment behavior. Models like o3-mini and Claude exhibit high confidence, often reporting over 80% coherence even in the zero-shot setting, suggesting strong internal certainty—though this may reflect overconfidence rather than accurate introspection. In contrast, models like GPT-4o and Llama show much more conservative estimates, with coherence below 30%, indicating either better-calibrated uncertainty or limited self-awareness. Interestingly, refinement tends to reduce overconfidence for some models (e.g., Claude) while slightly improving coherence estimation for others (e.g., GPT-4o and Deepseek), suggesting iterative attempts help align perceived and actual correctness.
We manually inspected a subset of successful solutions to validate if generated specifications align with the intended problem. Except for o3-mini, most models appear honest in their coherence self-assessments, with no false negatives found. Thus, our evaluation reflects an optimistic upper bound on model performance—assuming coherence estimates are accurate and verifier passes indicate best-case correctness. Automatically verifying the alignment between generated specifications and user intent in natural language remains an open technical challenge [24]. Tackling this alignment challenge is beyond the scope of this benchmark work and is left for future research.
Key Insights: Models show a wide range in coherence confidence level, suggesting various internal behaviors. On average, only 43% of passed solutions are judged coherent and our manual review suggests strong alignment.
4.6 Performance by Tools
Table 4: Average Pass Rates across Tools
Attempt CBMC Dafny Frama-C VerCors VeriFast Verus Why3
VerifyThisBench zero-shot 4.76% 1.30% 0.79% 2.31% 0 3.32% 0.51%
refinement 18.11% 4.47% 4.26% 5.34% 0.43% 8.15% 3.75%
VerifyThisBenchXS zero-shot - 1.35% 0.82% 0 5.22% - 0.21%
refinement - 21.47% 4.53% 1.28% 9.93% - 6.64%
Table 4 shows that all tools benefit from iterative refinement through feedback. In the VerifyThisBench setting, CBMC and Verus exhibit the most pronounced improvements, likely due to their syntactic similarity to C and Rust, making them more accessible to language models. Dafny also shows moderate gains in this setting.
In VerifyThisBenchXS, improvements are even more substantial. Dafny, in particular, demonstrates a remarkable leap—from near-zero success to over a 24% pass rate by Iteration 5, highlighting strong synergy with guided synthesis. In contrast, tools such as VeriFast, Frama-C, and Why3 remain largely stagnant on both benchmarks, suggesting either stricter syntactic or semantic constraints, or a structural mismatch with current model capabilities.
4.7 Performance by Relexation
Table 5: Overall Performance across Different Relaxation Settings
Code Specification Proof
Model Zero-shot Refinement Zero-shot Refinement Zero-shot Refinement
GPT4o 0.51% 10.26% 1.53% 8.67% 1.11% 4.44%
GPT4omini 0.51% 4.10% 1.53% 3.57% 0 2.22%
o3mini 0 7.18% 1.53% 5.10% 2.22% 7.78%
o4mini 0.51% 13.33% 2.04% 15.31% 1.11% 15.56%
Claude 1.03% 12.82% 2.55% 15.82% 0 7.78%
Gemini 1.03% 5.13% 0.51% 2.55% 1.11% 2.22%
Llama 0.51% 11.79% 1.02% 13.27% 0 5.56%
Deepseek 0.51% 16.41% 4.08% 17.86% 3.33% 11.11%
Qwen 0.51% 3.08% 1.02% 3.57% 0 0
Overall 0.57% 9.34% 1.76% 9.52% 0.99% 6.30%
Table 5 categorizes performance based on three relaxation types: specification, where the implementation and proof are given and the model fills in the spec; code, where the spec is provided and the model completes both the implementation and proof; and proof, where the full solution is available except for the loop invariant, which the model must supply. Across all categories, performance improves notably from Iter 1 (zero-shot) to Iter 5 (refinement), indicating that iterative refinement or feedback significantly aids verification success.
Among the three, specification relaxation yields the highest overall pass rates, suggesting that models are most effective when reasoning about what a program is supposed to do, given a working implementation and its proof context. Completing code implementation falls between the two, showing that models can sometimes generate plausible code. Completing loop invariant, arguably the most abstract and logically demanding task, results in lowest pass rates, though still showing solid gains with retries. This points to the inherent difficulty models face in understanding and completing partial proofs.
Key Insights: Generating the entire solution holistically (with overall pass rate of 9.34%) may not be more difficult than generating a specific one, e.g., loop invariant (with overall pass rate of 6.3%).
5 Conclusion
In this work, we introduce VerifyThisBench and VerifyThisBenchXS to evaluate the formal verification capabilities of large language models, systematically assessing their performance across a range of tools, tasks, and relaxation settings. Despite the use of SOTA models, results show generally poor performance, particularly in strict end-to-end settings that require complete formal reasoning without assistance. These findings highlight significant gaps in current models’ ability to generate semantically and logically correct solutions in formal domains.
This study also has several limitations. Our coherence metric relies on the model’s self-assessment, which may not reliably reflect true specification alignment. Additionally, our exploration of iteration is limited to feedback-based refinement; more sophisticated approaches,such as fine-tuning on formal verification tasks or incorporating external verification guidance, may yield further insights and improvements. Future work should explore these directions to better understand and advance the formal reasoning capabilities of language models.
Appendix A Composition of VerifyThisBenchXS
Table 6 presents the composition of VerifyThisBenchXS, summarizing the number of verification tasks by tool and category. It includes counts of implementations, specifications, and loop-related examples for five verification tools: Dafny, Frama-C, Vercors, Verifast, and Why3. In total, the benchmark comprises 481 tasks, with 195 implementations, 196 specifications, and 90 proof-related examples.
Table 6: Composition of VerifyThisBenchXS
Tool Implementaion Specification Proof Total
Dafny 28 25 21 74
FramaC 15 15 24 54
VerCors 8 8 10 26
VeriFast 26 31 9 66
Why3 118 117 26 261
Total 195 196 90 481
Appendix B VerifyThisBenchXS Data Source
Table 7 lists the sources of solutions used to construct VerifyThisBenchXS. It includes the year of publication, the name of the verification challenge, the verification tool used, and the authors or contributors of each solution.
Table 7: Solution used to generate VerifyThisBenchXS.
Year Challenge Name Tool Authors
2024 The Rope Data Structure Why3 Jean-Christophe Filliâtre
2024 Smart Array Copy by Shuffled Subsegments Why3 Jean-Christophe Filliâtre
2023 Binary Decision Diagrams Why3 Martin Clochard and Yannick Moy
2021 Lexicographic Permutations Why3 Jean-Christophe Filliâtre and Andrei Paskevich
2021 Lexicographic Permutations VerCors Marieke Huisman and Sebastiaan Joosten
2021 DLL to BST Why3 Jean-Christophe Filliâtre and Andrei Paskevich
2021 Shearsort Why3 Jean-Christophe Filliâtre and Andrei Paskevich
2019 Monotonic Segments and GHC sort Frama-C Virgile Prevosto and Virgile Robles
2019 Monotonic Segments and GHC sort Dafny Sample answer from report
2019 Cartesian Trees Frama-C Virgile Prevosto and Virgile Robles
2019 Sparse Matrix Multiplication Frama-C Virgile Prevosto and Virgile Robles
2018 Array Based Queuing Lock Why3 Raphael Rieu
2018 Gap buffer Why3 Raphael Rieu
2018 Colored tiles Why3 Raphael Rieu
2017 Pair Insertion Sort Frama-C Lionel Blatter and Jean-Christophe Léchenet
2017 Pair Insertion Sort Dafny Jon Mediero Iturrioz
2017 Pair Insertion Sort VerCors Marieke Huisman, Wytse Oortwijn
2017 Maximum-sum Array(one-dimension) Frama-C Lionel Blatter and Jean-Christophe Léchenet
2017 Odd-even Transposition Sort Frama-C Lionel Blatter and Jean-Christophe Léchenet
2017 Tree Buffer Frama-C Lionel Blatter and Jean-Christophe Léchenet
2017 Tree Buffer VerCors Marieke Huisman, Wytse Oortwijn
2016 Matrix Multiplication VeriFast Bart Jacobs
2016 Matrix Multiplication Dafny Luca Weibel and Christiaan Dirkx
2016 Matrix Multiplication Why3 Martin Clochard and Léon Gondelman and Mário Pereira
2016 Binary Tree Traversal VeriFast Bart Jacobs
2016 Binary Tree Traversal Why3 Martin Clochard and Léon Gondelman and Mário Pereira
2016 Static Tree Barrier VeriFast Bart Jacobs
2015 RELAXED PREFIX Why3 Jean-Christophe Filliâtre and Guillaume Melquiond
2015 PARALLEL GCD Why3 Jean-Christophe Filliâtre and Guillaume Melquiond
2015 DANCING LINKS Why3 Jean-Christophe Filliâtre and Guillaume Melquiond
2012 Longest Common Prefix VeriFast Bart Jacobs and Jan Smans
2012 Prefix-Sum VeriFast Bart Jacobs and Jan Smans
2012 Tree Del VeriFast Bart Jacobs and Jan Smans
2011 Finding the Maximum in an Array Dafny Julian Tschannen and Nadia Polikarpova
2011 Finding the Maximum in a Tree Dafny Julian Tschannen and Nadia Polikarpova
2011 Finding Two Duplets in an Array Dafny Julian Tschannen and Nadia Polikarpova
Appendix C Model Versions
GPT-4o was evaluated using the version from August 6, 2024, while GPT-4o-mini and o4-mini correspond to the July 18, 2024 versions. The o3-mini model was accessed as of January 31, 2025. Claude refers to the Claude-3.7-Sonnet version released on February 24, 2025, and Gemini-2.5 Flash on the April 17, 2025 release. For open-source models, we used LLaMA3.3-70b Instruct from December 6, 2024, DeepSeek-chat-v3 from March 24, 2025, and Qwen2.5-72b Instruct from September 19, 2024. These version references ensure the reproducibility and consistency of our benchmarking results.
Appendix D Tool Versions
We report the versions of all verification tools used in our experiments to ensure reproducibility. The Verus verifier was run using version v0.2025.04.03.0f22710, while Why3 was evaluated with version v1.6.0. For Frama-C, we used version v30.0, and VeriFast experiments were conducted with version v25.02. The Dafny toolchain ran on version v4.10.0, and Vercors with v2.3.0. Finally, we used CBMC version v6.5.0.
Below we briefly describe each tool:
  • Dafny: A verification-aware programming language with built-in specification support (pre/post-conditions, invariants) and an automatic static verifier.
  • Why3: A platform for deductive verification with its own intermediate language (WhyML) and integration with external theorem provers.
  • VeriFast: A verifier for C and Java using separation logic, enabling modular reasoning about memory safety and functional correctness.
  • VerCors: A verifier for concurrent programs in Java, C, and OpenCL, supporting permission-based separation logic and parallel reasoning.
  • Frama-C: A modular analysis platform for C, using the ACSL specification language and combining abstract interpretation with deductive verification.
  • Verus: A verifier for Rust programs that checks user-defined specifications using SMT solving, supporting low-level features and ownership semantics.
  • CBMC: A bounded model checker for C and C++ that verifies safety and functional correctness by translating code into SAT/SMT formulas.
Appendix E Example Challenge and Solution
1// # Description
2// This challenge is an instance of Kaldewaij’s Search by Elimination, where an element with a given property is located by eliminating elements that do not have thatproperty. The challenge was selected as it involves a relatively simple but interesting invariant, expressing that the maximal element is in the remaining search space rather than maintaining the maximal element found so far. A pseudo-code implementation is as follows:
3// int max(int[] a) {
4// int x = 0;
5// int y = a.length-1;
6// while (x != y) {
7// if (a[x] <= a[y]) x++;
8// else y--;
9// }
10// return x;
11// }
12// # Task
13// Please implementat and verify that the index returned by the method max() given points to an element maximal in the array
14pub fn max(a: &[int]) -> (result: usize)
15 requires
16 a.len() > 0,
17 ensures ({
18 result < a.len(),
19 forall|i: int| 0 <= i && i < a.len() ==> a[result as int] >= a[i],
20 }),
21{
22 let mut x: usize = 0;
23 let mut y: usize = a.len() - 1;
24 ...
25 while x != y
26 invariant
27 ...
28 x as int <= max_idx && max_idx <= y as int,
29 forall|i: int| 0 <= i && i < a.len() ==> a[max_idx] >= a[i],
30 decreases y - x
31 {
32 if a[x] <= a[y] {
34 proof {
35 ...
36 if max_idx == x as int {
37 assert(a[x] == a[y]);
38 max_idx = y as int;
39 }
40 ...
41 assert(x as int + 1 <= max_idx && max_idx <= y as int);
42 }
43 x += 1;
44 } else {
45 proof {
46 ...
47 }
48 y -= 1;
49 }
50 }
51 ...
52 return x;
53}
Figure 6: An example challenge stated in natural language highlighted in green and its potential solution in Verus with code implementation in grey, spec in yellow and proof in orange and invariants (a special kind of proof) in pink. This challenge is from 2011 and the solution is generated by Claude-3.7-Sonnet.